# Aeos: The Impact of Pop III Initial Mass Function and Star-by-Star Models in Galaxy Simulations

### v2

@@ 1 Introduction
The formation of the first stars, known as Population III (Pop III), marked a pivotal transition in the evolution of the early Universe (Klessen & Glover, 2023; Barkana & Loeb, 2001; Bromm, 2013; Bromm et al., 2002; Wise & Abel, 2008). Composed solely of primordial gas, these stars drove the chemical enrichment of the first galaxies. A key factor in determining the properties of these nascent galaxies is the initial mass function (IMF) of the Pop III stars (e.g., Schaerer, 2002; Schneider et al., 2006; Hartwig et al., 2015; Stacy et al., 2016; Ishigaki et al., 2018; Lazar & Bromm, 2022; Parsons et al., 2022). Understanding the Pop III IMF is crucial for constraining the feedback mechanisms that regulated star formation and the transition to later populations of stars.
@@
Early numerical simulations suggested that the Pop III IMF was top-heavy, favoring very massive stars reaching hundreds or even a thousand solar masses (Abel et al., 2002; Bromm et al., 2002). This result stemmed from the inability of the primordial gas to cool efficiently, leading to the formation of large single stars. However, more recent high-resolution simulations have suggested that gas cloud fragmentation could produce a broader range of stellar masses, extending down to tens of solar masses or even lower (Hosokawa et al., 2011; Hirano et al., 2014; Susa et al., 2014; Stacy et al., 2016). These results suggest that the Pop III IMF (which is still distinct from a classical Pop II IMF) encompasses a broad range of stellar masses, allowing for the formation of both extremely massive and relatively lower-mass stars in the early universe.
@@
Indirect observational evidence further supports the idea of a broad Pop III IMF. Metal-poor stars observed in the Milky Way and its satellite galaxies retain chemical abundance patterns that may represent enrichment by Pop III stars (Hartwig et al., 2018). When these patterns are compared to theoretical models of Pop III supernova yields, they can offer insights into the masses of Pop III progenitors. Some data suggest that certain enriched stars originated from Pop III progenitors with masses exceeding 50 M (Ji et al., 2024) and even up to 260 M (Xing et al., 2023); however, the latter finding is contested by recent work (Skúladóttir et al., 2024). Meanwhile, the majority of Pop III progenitors may have had masses of only a few tens of solar masses (Ishigaki et al., 2018; Fraser et al., 2017; Tumlinson, 2006). While some models propose a Salpeter-like IMF to describe the Pop III mass distribution, recent studies have pointed to possible deviations from this power law, including a potential peak around 25 M (Ishigaki et al., 2018). The non-detection of zero-metallicity stars in ultra-faint dwarf galaxies also provides constraints on the minimum mass of the first stars, $M_{\text{min}}>0.8$ M (Rossi et al., 2021).
@@
Despite these advances, significant uncertainties remain regarding the true shape of the Pop III IMF and its implications for galaxy formation and reionization. Simulations provide a valuable tool for exploring these uncertainties by modeling different IMF scenarios and examining their effects on key cosmic processes. The amount of ionizing radiation produced by Pop III stars, for instance, depends heavily on their mass distribution, directly impacting the ionization history of the universe and star formation in early galaxies (Schaerer, 2002; Wise et al., 2012a; Kimm et al., 2017). Likewise, stellar feedback from SNe and winds influences the growth and chemical evolution of the earliest galaxies (Wise et al., 2012b; Heger & Woosley, 2010).
@@
The legacy of the Pop III stars extends to the formation of Population II (Pop II) stars (e.g., Prgomet et al., 2022). The Pop III IMF influences the characteristics of the Pop II stars, which emerged after the interstellar medium was enriched by metals from Pop III SNe. Massive Pop II stars play a critical role in early galaxy evolution as they contribute significantly to the regulation of star formation through feedback processes such as ionizing radiation, stellar winds, and SNe (Somerville & Davé, 2015). These feedback mechanisms drive gas outflows, influence the ionization history, and regulate the growth of early galaxies. Accurately capturing the effects of individual Pop II stars on their environment is therefore also essential for understanding the formation and evolution of the first galaxies (e.g., Brauer et al., 2025; Andersson et al., 2025).
@@
Traditional cosmological simulations often simplify the modeling of Pop II stellar feedback by representing entire star clusters as single particles, which limits the resolution of the feedback processes that individual stars provide (e.g., Skinner & Wise, 2020; Smith, 2021). Such models can overlook the detailed radiative and mechanical feedback from individual stars, which plays a critical role in regulating the interstellar medium and shaping galaxy evolution. In recent years, simulations have begun to include feedback from individual stars in galaxy simulations (e.g., Emerick et al., 2019; Lahén et al., 2020; Hirai et al., 2021; Gutcke et al., 2021; Hislop et al., 2022; Andersson et al., 2023; Calura et al., 2022; Steinwandel et al., 2023; Brauer et al., 2025; Hirai et al., 2024; Deng et al., 2024). Star-by-star modeling of Pop II stellar feedback offers a more detailed approach. This captures the impact of the radiation, supernovae (SNe), and stellar winds from individual stars both inside and outside clusters on gas dynamics and chemical enrichment, which is crucial for understanding the earliest stages of galaxy formation.
@@
In this work, we present results from the Aeos simulations (Brauer et al., 2025; Mead et al., 2025) that explore (1) the impact of varying the Pop III IMF on early galaxy evolution and (2) the impact of star-by-star models when simulating the first galaxies. These Aeos simulations model the formation of the first stars and galaxies in the first few hundred megayears of the universe, following individual stellar feedback and detailed chemical enrichment. This approach is motivated by the fundamental importance of predicting the chemical abundances of ancient stars, particularly those found in ultra-faint dwarf galaxies and the Milky Way’s stellar halo, to unravel the processes of early star and galaxy formation (e.g., Brauer et al., 2019).
@@ 2 Methods
This study uses the numerical methods detailed in Brauer et al. (2025). We focus on a comparative analysis between the fiducial Aeos cosmological simulation and additional simulations: (1) an Aeos simulation with an altered Pop III IMF with a higher characteristic mass $M_{\rm char}$ and upper mass cutoff, and (2) seven comparison simulations that omit the full star-by-star feedback and enrichment models, instead using single stellar population particles and IMF averaged metal yields. The comparison simulations are identical except for differences arising primarily from stochastic IMF sampling during star formation. We run seven comparison simulations to show how stochasticity affects results. Below, we provide a concise summary of the important methods from Brauer et al. (2025) and elaborate on the modifications pertinent to the additional simulations. The differences in the simulations are summarized in Table 1.
@@ 2.1 Similarities Across Simulations
All simulations use the Enzo adaptive mesh refinement (AMR) code (Bryan et al., 2014; Brummel-Smith et al., 2019), enabling high-resolution modeling of cosmological structures. The computational grid employs AMR to achieve high physical resolution in regions of elevated gas density, ensuring detailed resolution of dense gas clumps and star formation. They have a root-grid resolution of $256^{3}$, a dark matter resolution of 1840 M, and a resolution of 1 physical pc at the finest scales.
The simulations cover a comoving (1 Mpc)3 volume, simulated from redshift $z=130$ to $z=14.5$, approximately 300 Myr after the Big Bang. They commence from identical initial conditions generated with MUSIC (Hahn & Abel, 2011) at redshift $z=130$, adhering to the cosmological parameters derived from the Planck collaboration (Planck Collaboration et al., 2014). This ensures consistency across all simulation runs, facilitating comparisons between different feedback models and IMF choices.
Star formation is modeled stochastically in cold, dense gas regions with $n>10^{4}$ cm-3 that exhibit converging flows having $\nabla\cdot\mathbf{v}<0$, assuming a star formation efficiency per free-fall time of $e_{\rm ff}=2\%$. The star formation algorithm distinguishes Pop II from Pop III stars using a threshold of total gas metallicity $Z>10^{-5}Z_{\odot}$ (Ji et al., 2014; Chiaki et al., 2015; Schneider et al., 2012; Tumlinson, 2006). Pop III star formation in lower metallicity gas further requires a molecular hydrogen fraction $f_{\rm H_{2}}>0.0005$, consistent with high-resolution simulations (Susa et al., 2014; Kulkarni et al., 2021).
Following Wise et al. (2012b), we adopt a Pop III IMF of the form:
@@
Above $M_{\text{char}}$, it behaves as a Salpeter IMF, and below $M_{\text{char}}$ it is exponentially suppressed (Chabrier, 2003). We also include an upper mass limit $M_{\rm max}$. In all simulations, every Pop III star is represented by a single star particle. The Pop II IMF is sampled from a Kroupa (2001) distribution with a mass range of 0.08–120 M.
@@
We follow radiation from every star with mass $M>8$ M. For Pop III stars, we use binned photon counts from Heger & Woosley (2010) with the lifetimes in Schaerer (2002). For Pop II stars, photon fluxes in each radiation band are determined using the OSTAR2002 (Lanz & Hubeny, 2003) grid of O-type stellar models. We also use the PARSEC (Bressan et al., 2012) grid of stellar evolution tracks to set the lifetime of each Pop II star and the start time and length of the asymptotic giant branch (AGB) phase, if present.
@@ 2.2 Differences Between Simulations
Aeos10 adopts a Pop III IMF with a characteristic mass $M_{\rm char}=10$ M and a maximum mass of $M_{\rm max}=100$ M. To investigate the sensitivity of galaxy evolution to the Pop III IMF, we execute an additional Aeos simulation (Aeos20) wherein the characteristic mass $M_{\rm char}$ of the Pop III IMF is modified from the fiducial value of 10 M to 20 M and $M_{\rm max}=300$ M (motivated by Hirano & Bromm, 2017; Bromm, 2013; Yoshida, 2006; Skinner & Wise, 2020). This adjustment alters the mass distribution of Pop III stars, thereby affecting their lifetimes, feedback outputs, and chemical yields (see Brauer et al., 2025). This is the only difference between the Aeos10 and Aeos20 simulations.
@@
Additionally, the comparison simulations track overall metallicity but do not track individual elements. The Aeos simulations track the evolution of ten individual metal abundances – C, N, O, Na, Mg, Ca, Mn, Fe, Sr, and Ba – in addition to hydrogen and helium. These elements are sourced from various nucleosynthetic channels, including core-collapse SNe, Type Ia SNe, and asymptotic giant branch stars. We use detailed yield tables from Heger & Woosley (2010) for Pop III CCSNe and Limongi & Chieffi (2018) for Pop II stars.
@@ 3.1 Ionization
This occurs because the ionizing radiation emitted by a star increases with its mass (Schaerer, 2002). Consequently, the IMF in Aeos20, with its higher $M_{\text{char}}$ and upper mass limit, produces a stellar population that emits significantly more ionizing photons. Figure 2 shows the mass distribution of Pop III stars in the two simulations. Aeos10 contains approximately 2000 stars with masses in the tens of solar mass range, while Aeos20 has fewer than 500 stars in this range. However, Aeos20 features Pop III stars up to 300 M, including about 80 stars exceeding 100 M. The total number of ionizing photons from Pop III stars per stellar mass for these two distributions (combining the histograms and rate of ionizing photons shown in Figure 2) is about $7\times 10^{61}$ photons/M for Aeos10 and $10\times 10^{61}$ photons/M for Aeos20.
@@
The higher ionization in Aeos20 suppresses the formation of small galaxies. Previous work has found that ionization-driven photoheating of the gas in and around dwarf galaxies suppresses star formation (Efstathiou, 1992; Bullock et al., 2000), particularly in halos with $M_{\text{halo}}\lesssim 10^{9}$ M (e.g., Thoul & Weinberg, 1996; Okamoto et al., 2008; Dawoodbhoy et al., 2018). In our simulations, all galaxies exist in small halos and we find they are highly sensitive to the rate of reionization. As seen in Figure 4, Aeos10 has far more galaxies at low halo masses. For $M_{*}<100$ M, Aeos10 has 60 star-forming halos while Aeos20 has only 14 at $z=14.5$. Meanwhile, the distributions of halo masses for all halos are essentially identical for Aeos10, Aeos20, and the comparison simulations, as expected for simulations with the same dark matter initialization. Another mechanism, Lyman-Werner (LW) radiation, can also suppress early star formation by dissociating H2, the primary coolant in metal-poor gas (e.g., Safranek-Shrader et al., 2012). Both Aeos10 and Aeos20 have identical LW backgrounds, however, so the suppression of small galaxy formation in Aeos20 vs. Aeos10 is isolated to the differences in ionizing photons. At the beginning of the simulation ($\sim 120$ Myr), Aeos10 briefly surpasses Aeos20 in ionizing photons (see Figures 1 and 3). This is due to stochastic differences during the beginning of Pop III star formation, when there are very few stars, and a slightly delayed onset of Pop III star formation in Aeos20. An ensemble of simulations with Aeos10 and Aeos20 parameters would be preferred to better understand stochastic differences, but is currently unfeasible due to the computational cost of the simulations.
We selected Aeos10 as our fiducial simulation in Brauer et al. (2025) because the high level of ionization in Aeos20 causes reionization to occur much earlier than expected: 30% ionized at redshift $z\sim 14$ compared to the inferred redshift of $z\sim 8$ from Planck Collaboration et al. (2016). Our simulation volume is a small region that is not generally representative of the very early Universe, however. We thus avoid interpreting this early reionization as a constraint on the Pop III IMF, which remains highly uncertain. Instead, we interpret the differences between Aeos10 and Aeos20 as evidence of the strong impact that different Pop III IMF choices have on reionization and the number of small, early galaxies. Given both the importance of the Pop III IMF and its uncertainty, simulations of the first galaxies must take care to understand the impact of Pop III IMF choices on their results.

### v3

# Aeos: The Impact of Population III Initial Mass Function and Star-by-Star Models in Galaxy Simulations

@@ 1 Introduction
The formation of the first stars, known as Population III (Pop III), marked a pivotal transition in the evolution of the early Universe (Klessen2023; Barkana2001; Bromm13; Bromm2002; Wise2008). Composed solely of primordial gas, these stars drove the chemical enrichment of the first galaxies. A key factor in determining the properties of these nascent galaxies is the initial mass function (IMF) of the Pop III stars (e.g., Schaerer2002; Schneider2006; Hartwig2015; Stacy2016; Ishigaki2018; Lazar2022; Parsons2022). Understanding the Pop III IMF is crucial for constraining the feedback mechanisms that regulated star formation and the transition to later populations of stars.
@@
Early numerical simulations suggested that the Pop III IMF was top-heavy, favoring very massive stars reaching hundreds or even a thousand solar masses (Abel2002; Bromm2002). This result stemmed from the inability of the primordial gas to cool efficiently, leading to the formation of large single stars. However, more recent high-resolution simulations have suggested that gas cloud fragmentation could produce a broader range of stellar masses, extending down to tens of solar masses or even lower (Hosokawa11; Hirano14; Susa2014; Stacy2016). These results suggest that the Pop III IMF (which is still distinct from a classical Pop II IMF) encompasses a broad range of stellar masses, allowing for the formation of both extremely massive and relatively lower-mass stars in the early universe.
@@
Indirect observational evidence further supports the idea of a broad Pop III IMF. Metal-poor stars observed in the Milky Way and its satellite galaxies retain chemical abundance patterns that may represent enrichment by Pop III stars (Hartwig2018). When these patterns are compared to theoretical models of Pop III supernova yields, they can offer insights into the masses of Pop III progenitors. Some data suggest that certain enriched stars originated from Pop III progenitors with masses exceeding 50 M (Ji24) and even up to 260 M (Xing23); however, the latter finding is contested by recent work (Skuladottir2024). Meanwhile, the majority of Pop III progenitors may have had masses of only a few tens of solar masses (Ishigaki2018; Fraser17; Tumlinson2006). While some models propose a Salpeter-like IMF to describe the Pop III mass distribution, recent studies have pointed to possible deviations from this power law, including a potential peak around 25 M (Ishigaki2018). The non-detection of zero-metallicity stars in ultra-faint dwarf galaxies also provides constraints on the minimum mass of the first stars, $M_{\text{min}}>0.8$ M (Rossi21).
@@
Despite these advances, significant uncertainties remain regarding the true shape of the Pop III IMF and its implications for galaxy formation and reionization. Simulations provide a valuable tool for exploring these uncertainties by modeling different IMF scenarios and examining their effects on key cosmic processes. The amount of ionizing radiation produced by Pop III stars, for instance, depends heavily on their mass distribution, directly impacting the ionization history of the universe and star formation in early galaxies (Schaerer2002; WiseAbel2012; Kimm2017). Likewise, stellar feedback from SNe and winds influences the growth and chemical evolution of the earliest galaxies (Wise2012a; HegerWoosley2010).
@@
The legacy of the Pop III stars extends to the formation of Population II (Pop II) stars (e.g., Prgomet22). The Pop III IMF influences the characteristics of the Pop II stars, which emerged after the interstellar medium was enriched by metals from Pop III SNe. Massive Pop II stars play a critical role in early galaxy evolution as they contribute significantly to the regulation of star formation through feedback processes such as ionizing radiation, stellar winds, and SNe (SomervilleDave2015). These feedback mechanisms drive gas outflows, influence the ionization history, and regulate the growth of early galaxies. Accurately capturing the effects of individual Pop II stars on their environment is therefore also essential for understanding the formation and evolution of the first galaxies (e.g., AeosMethods; Andersson24).
@@
Traditional cosmological simulations often simplify the modeling of Pop II stellar feedback by representing entire star clusters as single particles, which limits the resolution of the feedback processes that individual stars provide (e.g., Skinner20; Smith2021). Such models can overlook the detailed radiative and mechanical feedback from individual stars, which plays a critical role in regulating the interstellar medium and shaping galaxy evolution. In recent years, simulations have begun to include feedback from individual stars in galaxy simulations (e.g., Emerick2019; Lahen2020; Hirai21; Gutcke21; Hislop22; Andersson23; Calura22; Steinwandel23; AeosMethods; Hirai24; Deng24). Star-by-star modeling of Pop II stellar feedback offers a more detailed approach. This captures the impact of the radiation, supernovae (SNe), and stellar winds from individual stars both inside and outside clusters on gas dynamics and chemical enrichment, which is crucial for understanding the earliest stages of galaxy formation.
@@
In this work, we present results from the Aeos simulations (AeosMethods; Mead2024b) that explore (1) the impact of varying the Pop III IMF on early galaxy evolution and (2) the impact of star-by-star models when simulating the first galaxies. These Aeos simulations model the formation of the first stars and galaxies in the first few hundred megayears of the universe, following individual stellar feedback and detailed chemical enrichment. This approach is motivated by the fundamental importance of predicting the chemical abundances of ancient stars, particularly those found in ultra-faint dwarf galaxies and the Milky Way’s stellar halo, to unravel the processes of early star and galaxy formation (e.g., Brauer19).
@@ 2 Methods
This study uses the numerical methods detailed in AeosMethods. We focus on a comparative analysis between the fiducial Aeos cosmological simulation and additional simulations: (1) an Aeos simulation with an altered Pop III IMF with a higher characteristic mass $M_{\rm char}$ and upper mass cutoff, and (2) seven comparison simulations that omit the full star-by-star feedback and enrichment models, instead using single stellar population particles and IMF averaged metal yields. The comparison simulations are identical except for differences arising primarily from stochastic IMF sampling during star formation. We run seven comparison simulations to show how stochasticity affects results. Below, we provide a concise summary of the important methods from AeosMethods and elaborate on the modifications pertinent to the additional simulations. The differences in the simulations are summarized in Table 1.
@@ 2.1 Similarities Across Simulations
All simulations use the Enzo adaptive mesh refinement (AMR) code (Enzo2014; Enzo2019), enabling high-resolution modeling of cosmological structures. The computational grid employs AMR to achieve high physical resolution in regions of elevated gas density, ensuring detailed resolution of dense gas clumps and star formation. They have a root-grid resolution of $256^{3}$, a dark matter resolution of 1840 M, and a resolution of 1 physical pc at the finest scales.
The simulations cover a comoving (1 Mpc)3 volume, simulated from redshift $z=130$ to $z=14.5$, approximately 300 Myr after the Big Bang. They commence from identical initial conditions generated with MUSIC (Hahn2011) at redshift $z=130$, adhering to the cosmological parameters derived from the Planck collaboration (Planck2014). This ensures consistency across all simulation runs, facilitating comparisons between different feedback models and IMF choices.
Star formation is modeled stochastically in cold, dense gas regions with $n>10^{4}$ cm-3 that exhibit converging flows having $\nabla\cdot\mathbf{v}<0$, assuming a star formation efficiency per free-fall time of $e_{\rm ff}=2\%$. The star formation algorithm distinguishes Pop II from Pop III stars using a threshold of total gas metallicity $Z>10^{-5}Z_{\odot}$ (Ji2014; Chiaki2015; Schneider2012; Tumlinson2006). Pop III star formation in lower metallicity gas further requires a molecular hydrogen fraction $f_{\rm H_{2}}>0.0005$, consistent with high-resolution simulations (Susa2014; Kulkarni2021).
Following Wise2012a, we adopt a Pop III IMF of the form:
@@
Above $M_{\text{char}}$, it behaves as a Salpeter IMF, and below $M_{\text{char}}$ it is exponentially suppressed (Chabrier2003). We also include an upper mass limit $M_{\rm max}$. In all simulations, every Pop III star is represented by a single star particle. The Pop II IMF is sampled from a Kroupa2001 distribution with a mass range of 0.08–120 M.
@@
We follow radiation from every star with mass $M>8$ M. For Pop III stars, we use binned photon counts from HegerWoosley2010 with the lifetimes in Schaerer2002. For Pop II stars, photon fluxes in each radiation band are determined using the OSTAR2002 (Lanz2003) grid of O-type stellar models. We also use the PARSEC (Bressan2012) grid of stellar evolution tracks to set the lifetime of each Pop II star and the start time and length of the asymptotic giant branch (AGB) phase, if present.
@@ 2.2 Differences Between Simulations
Aeos10 adopts a Pop III IMF with a characteristic mass $M_{\rm char}=10$ M and a maximum mass of $M_{\rm max}=100$ M. To investigate the sensitivity of galaxy evolution to the Pop III IMF, we execute an additional Aeos simulation (Aeos20) wherein the characteristic mass $M_{\rm char}$ of the Pop III IMF is modified from the fiducial value of 10 M to 20 M and $M_{\rm max}=300$ M (motivated by Hirano17; Bromm13; Yoshida06; Skinner20). This adjustment alters the mass distribution of Pop III stars, thereby affecting their lifetimes, feedback outputs, and chemical yields (see AeosMethods). This is the only difference between the Aeos10 and Aeos20 simulations.
@@
Additionally, the comparison simulations track overall metallicity but do not track individual elements. The Aeos simulations track the evolution of ten individual metal abundances – C, N, O, Na, Mg, Ca, Mn, Fe, Sr, and Ba – in addition to hydrogen and helium. These elements are sourced from various nucleosynthetic channels, including core-collapse SNe, Type Ia SNe, and asymptotic giant branch stars. We use detailed yield tables from HegerWoosley2010 for Pop III CCSNe and Limongi2018 for Pop II stars.
@@ 3.1 Ionization
This occurs because the ionizing radiation emitted by a star increases with its mass (Schaerer2002). Consequently, the IMF in Aeos20, with its higher $M_{\text{char}}$ and upper mass limit, produces a stellar population that emits significantly more ionizing photons. Figure 2 shows the mass distribution of Pop III stars in the two simulations. Aeos10 contains approximately 2000 stars with masses in the tens of solar mass range, while Aeos20 has fewer than 500 stars in this range. However, Aeos20 features Pop III stars up to 300 M, including about 80 stars exceeding 100 M. The total number of ionizing photons from Pop III stars per stellar mass for these two distributions (combining the histograms and rate of ionizing photons shown in Figure 2) is about $7\times 10^{61}$ photons/M for Aeos10 and $10\times 10^{61}$ photons/M for Aeos20.
@@
The higher ionization in Aeos20 suppresses the formation of small galaxies. Previous work has found that ionization-driven photoheating of the gas in and around dwarf galaxies suppresses star formation (Efstathiou1992; Bullock2000), particularly in halos with $M_{\text{halo}}\lesssim 10^{9}$ M (e.g., Thoul1996; Okamoto2008; Dawoodbhoy2018). In our simulations, all galaxies exist in small halos and we find they are highly sensitive to the rate of reionization. As seen in Figure 4, Aeos10 has far more galaxies at low halo masses. For $M_{*}<100$ M, Aeos10 has 60 star-forming halos while Aeos20 has only 14 at $z=14.5$. Meanwhile, the distributions of halo masses for all halos are essentially identical for Aeos10, Aeos20, and the comparison simulations, as expected for simulations with the same dark matter initialization. Another mechanism, Lyman-Werner (LW) radiation, can also suppress early star formation by dissociating H2, the primary coolant in metal-poor gas (e.g., Safranek2012). Both Aeos10 and Aeos20 have identical LW backgrounds, however, so the suppression of small galaxy formation in Aeos20 vs. Aeos10 is isolated to the differences in ionizing photons. At the beginning of the simulation ($\sim 120$ Myr), Aeos10 briefly surpasses Aeos20 in ionizing photons (see Figures 1 and 3). This is due to stochastic differences during the beginning of Pop III star formation, when there are very few stars, and a slightly delayed onset of Pop III star formation in Aeos20. An ensemble of simulations with Aeos10 and Aeos20 parameters would be preferred to better understand stochastic differences, but is currently unfeasible due to the computational cost of the simulations.
We selected Aeos10 as our fiducial simulation in AeosMethods because the high level of ionization in Aeos20 causes reionization to occur much earlier than expected: 30% ionized at redshift $z\sim 14$ compared to the inferred redshift of $z\sim 8$ from planck2016. Our simulation volume is a small region that is not generally representative of the very early Universe, however. We thus avoid interpreting this early reionization as a constraint on the Pop III IMF, which remains highly uncertain. Instead, we interpret the differences between Aeos10 and Aeos20 as evidence of the strong impact that different Pop III IMF choices have on reionization and the number of small, early galaxies. Given both the importance of the Pop III IMF and its uncertainty, simulations of the first galaxies must take care to understand the impact of Pop III IMF choices on their results.